OpenStudy (anonymous):

@terenzreignz how do I find the slope of a line that is parallel and perpendicular to y=4x-8

terenzreignz (terenzreignz):

Parallel lines have the same slope. So the slope of a line parallel to this one would be...?

OpenStudy (anonymous):

4

terenzreignz (terenzreignz):

Therefore, the slope of a line parallel to this one would also be 4. Perpendicular lines are tricky, what you need to remember is that if two lines are perpendicular, then the product of their slopes would be -1. So if m is the slope of a line perpendicular to this one, then just solve 4m = -1

OpenStudy (anonymous):

-5

terenzreignz (terenzreignz):

4m = -1 NOT 4 + m = -1 There's a difference. ;) 4m = -1 please solve for m (again)

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

right?

terenzreignz (terenzreignz):

No. that'd be correct if it were 4 + m = -1 instead. but it's 4m = -1 4 and m are multiplied, not added. What you do is divide both sides by 4.

OpenStudy (anonymous):

m=-1/4

terenzreignz (terenzreignz):

better :) and that's the slope of a line perpendicular to your line.

OpenStudy (anonymous):

Thanks so much @terenzreignz

terenzreignz (terenzreignz):

Call me TJ. And no problem ;)

OpenStudy (anonymous):

I have another question similar to this one, 3x+y=3(x+2) can we both solve it and compare answers? thanks TJ :)